# Tight concentration inequalities for quantum adversarial setups exploiting permutation symmetry

### v2

@@ III Numerical comparison of the performance with other concentration inequalities
III.1 Toy model for statistical estimation
@@
$\begin{split}{\cal A}(\gamma,\Delta)&\coloneqq\bigl{\{}(p_{1},p_{2},p_{3}):p_{1},p_{2},p_{3}\geq 0,\,p_{1}+p_{2}+p_{3}=1,\\ &\hskip 75.39963ptp_{1}-\gamma p_{2}\geq m(\gamma)+\Delta\bigr{\}},\end{split}$ (150)
@@
$M_{c}(\gamma)\leq m(\gamma)I,\qquad m(\gamma)\coloneqq\frac{1-\gamma+\sqrt{(1-\gamma)^{2}+4r\gamma}}{4},$ (152)
@@
$\mathrm{Pr}\left[\frac{1}{N}\sum_{i}\left(\hat{c}_{i}-\mathbb{E}[\hat{c}_{i}\mid\hat{c}_{<i}]\right)\geq t\right]\leq\exp\bigl{[}-2Nt^{2}/(1+\gamma)^{2}\bigr{]},$ (157)
@@
$\mathrm{Pr}\left[\frac{\hat{A}}{N}-\gamma\frac{\hat{B}}{N}\geq m(\gamma)+\Delta\right]\leq\exp\bigl{[}-2N\Delta^{2}/(1+\gamma)^{2}\bigr{]}.$ (160)
@@
$\epsilon_{\rm azuma}(\gamma,N,\Delta)=\exp\bigl{[}-2Ng^{2}/(1+\gamma)^{2}\bigr{]}$ (162)
@@
$\begin{split}&\mathrm{Pr}\biggl{[}\sum_{i=1}^{N}\hat{a}_{i}-\mathbb{E}[\hat{a}_{i}\mid\hat{a}_{<i},\hat{b}_{<i}]-\gamma\left(\hat{b}_{i}-\mathbb{E}[\hat{b}_{i}\mid\hat{a}_{<i},\hat{b}_{<i}]\right)\\ &\qquad\geq\gamma\sqrt{N}\biggl{(}\beta^{*}-\alpha^{*}+\frac{2\alpha^{*}}{N}\sum_{i=1}^{N}\hat{b}_{i}\biggr{)}+\sqrt{\frac{N\ln(2/\epsilon)}{2}}\biggr{]}\\ &\hskip 199.16928pt\leq\epsilon,\end{split}$ (166)
@@
$\begin{split}\Delta_{\rm kato}(\gamma,N,\epsilon)=\frac{\gamma}{\sqrt{N}}\biggl{[}\beta^{*}-\alpha^{*}\biggl{(}1-\frac{2\hat{B}}{N}\biggr{)}\biggr{]}+\sqrt{\frac{\ln(2/\epsilon)}{2N}}\end{split}$ (169)
achieves Eq. (154) in this case. Due to the use of the union bound, the derivation of a concentration inequality of the type Eq. (155) in this case requires additional optimization on how to assign the deviation $\Delta$ for $\hat{a}_{i}$ and $\hat{b}_{i}$. To avoid the complication, we do not consider a concentration inequality of the type Eq. (155) for Kato’s inequality.
@@
Now, assume that the i.i.d. state $\rho(\bm{q})^{\otimes n}$ is prepared with $\rho(\bm{q})$ giving the probability mass function $\bm{q}\in{\cal Q}$. Then, the probability that the sums of the outcomes of these measurements $\hat{A}$ and $\hat{B}$ lies in the region ${\cal A}(\gamma,\Delta_{q})$ is given from the refined version of the Sanov’s theorem [2, 3] as
$\begin{split}&\mathrm{Pr}\left[\left(\frac{\hat{A}}{N},\frac{\hat{B}}{N},1-\frac{\hat{A}}{N}-\frac{\hat{B}}{N}\right)\in{\cal A}(\gamma,\Delta_{q})\middle|\rho(\bm{q})^{\otimes n}\right]\\ &\leq\max_{\bm{p}\in{\cal A}(\gamma,\Delta_{q})}\exp[-ND(\bm{p}\|\bm{q})].\end{split}$ (171)
@@
$\begin{split}&\mathrm{Pr}\left[\left(\frac{\hat{A}}{N},\frac{\hat{B}}{N},1-\frac{\hat{A}}{N}-\frac{\hat{B}}{N}\right)\in{\cal A}(\gamma,\Delta_{q})\middle|\rho^{\otimes N}\right]\\ &\leq\max_{\bm{p}\in{\cal A}(\gamma,\Delta_{q})}\max_{\bm{q}\in{\cal Q}}\exp[-ND(\bm{p}\|\bm{q})].\end{split}$ (172)
Now, from Theorem 1, we have, for any permutation-invariant state $\rho_{\rm sym}$,
@@
then we obtain a concentration inequality of the type Eq. (155).
We finally consider applying Theorem 3 to this problem. As previously mentioned, an independent and identical measurement is performed on an unknown quantum state and the set of measurement outcomes is given by $\{(\text{I},0),(\text{I},1),(\text{I\!I},0),(\text{I\!I},1)\}$. Since we do not distinguish $(\text{I\!I},0)$ and $(\text{I\!I},1)$ as mentioned previously, we regard this measurement as three-outcome measurement and set $k=3$. For the set ${\cal Q}$ of probability mass functions that can be realized by this measurement given in Eq. (170), we consider the following set ${\cal R}(\gamma^{\prime})\supseteq{\cal Q}$ whose boundary is given by a hyperplane $R(\gamma^{\prime})$:
@@
The fact that this includes ${\cal Q}$ as a subset follows from Eqs. (151) and (152). Due to the requirement of Theorem 3 that the complement ${\cal P}\setminus{\cal R}(\gamma^{\prime})$ of ${\cal R}(\gamma^{\prime})$ contains only one probability mass function that corresponds to having a deterministic outcome, the parameter $\gamma^{\prime}$ is chosen to satisfy $10\leq m(\gamma^{\prime})<$ 1. Likewise, the region ${\cal R}^{\prime}(\gamma^{\prime})$ can be defined through Eqs. (132) and (55) by
@@
Figure 2: Comparison of the deviation $\Delta(\gamma,N,\epsilon)$ from (an upper bound on) the expectation value in Eqs. (150) and (154) when allowing the failure probability $\epsilon$. The parameters are chosen to be $\epsilon=10^{-30}$, $r=0.01$, and $B^{*}=0.01N$, and $\gamma$ is optimized to be $1.662$ so that an unconfirmed upper bound $\gamma B^{*}/N+m(\gamma)$ on $\hat{A}/N$ is minimized to be $0.02873$. We further assume that the random variable $\hat{B}$ is equal to the expected value $B^{*}$. The functions $\Delta_{\rm azuma}$ and $\Delta_{\rm kato}$ are defined respectively in Eqs. (161) and (169), and the functions $\Delta_{q}$, and $\Delta_{c}$ are defined implicitly through Eqs. (173) and (177).
@@
We numerically demonstrate the tightness of these bounds in some cases. The first example is the case that is relevant to the finite-size analysis of the QKD. In QKD, the parameter $\epsilon$ is fixed and its value is typically chosen to be $\sim 10^{-30}$. In the QKD application, we have the expected value $B^{*}$ of $\hat{B}$ from the channel that we use for the transmission, which is the reason why Kato’s inequality [12] typically gives a tighter bound in the finite-size regime. For a model, we set $\epsilon=10^{-30}$, $r=0.01$, and $B^{*}=0.01N$. The parameter $\gamma$ in ${\cal A}(\gamma,\Delta)$ is chosen so that the expected (unconfirmed) upper bound $\gamma B^{*}/N+m(\gamma)$ on $\hat{A}/N$ in the limit $N\rightarrow\infty$ is minimized. In the choice of parameters above, we have $\gamma=1.662$ to give the expected asymptotic upper bound $0.02873$ on $\hat{A}/N$. We further assume the case in which $\hat{B}$ is exactly equal to $B^{*}$. Under these assumptions, we derive the deviation from the above asymptotic upper bound for finite $N$ allowing the failure probability $\epsilon$ with the four methods described above. Figure 2 shows the deviation $\Delta(\gamma,N,\epsilon)$ when varying the number $N$ of quantum systems to be measured. As the figure suggests, $\Delta_{c}$ defined through Eq. 177 is the smallest among four bounds in this setup and the above particular choices of parameters except for some points that may be caused by the insufficient optimization. We can also see that $\Delta_{q}$ defined through Eq. (173) gives the comparable bound. They are smaller than $\Delta_{\rm kato}$ defined in Eq. (169), while $\Delta_{\rm kato}$ is still smaller than the $\Delta_{\rm azuma}$ defined in Eq. (161).
@@
Figure 3: Comparison of the minus logarithm of the failure probability $\epsilon(\gamma,N,\Delta)$ in Eq. (155) against the number $N$ of measured quantum systems for a fixed value of the deviation $\Delta$. The parameters are chosen to be $\epsilon=10^{-30}$, $r=0.01$, and $B^{*}=0.01N$, and $\gamma$ is optimized to be $1.662$ so that an unconfirmed upper bound $\gamma B^{*}/N+m(\gamma)$ on $\hat{A}/N$ is minimized to be $0.02873$. We further assume that the random variable $\hat{B}$ is equal to the expected value $B^{*}$. The functions $\epsilon_{\rm azuma}$, $\epsilon_{q}$, and $\epsilon_{c}$ are defined in Eqs. (162), (174), and (178), respectively.
@@
The next example is the case when we fix the deviation $\Delta$ from the expectation value (that is, an upper bound that holds in the asymptotic limit as shown in Eq. (158)) to a small value, and see how quickly the failure rate $\epsilon(\gamma,N,\Delta)$ in Eq. (155) goes to zero. We set the deviation $\Delta=0.01$, $r=0.01$ and $B^{*}=0.01N$, and choose $\gamma=1.662$ to give the asymptotically optimal upper bound $0.02873$ on $\hat{A}/N$ as was done in the previous figure. We also consider the case in which the observed value $\hat{B}$ is exactly equal to $B^{*}$. With these conditions, we plot $\epsilon_{\rm azuma}$ in Eq. (162), $\epsilon_{q}$ in Eq. (174), and $\epsilon_{c}$ in Eq. (178) in Fig. 3. As the figure suggests, the minus of each exponent, which governs the convergence speed, grows almost linearly for $N$, but its coefficient is very different between Azuma’s and ours. In fact, the convergence speed of our bounds is much faster than that of Azuma’s. Furthermore, $-\ln\epsilon_{q}$ and $-\ln\epsilon_{c}$ appear to have an almost constant gap. This gap exactly comes from the difference of the factors $f_{q}(N,d=2)\sim N^{\frac{3}{2}}$ and $f_{c}(N,k=3)\sim N$. We should note that there are regions in which Azuma’s inequality gives a better bound than ours when $N$ is small and the failure probability is large. When we alter the parameters $r$, $M$, and $\Delta$ in our model, then this region varies. However, our bound overall gives a faster convergence to the expectation value.
III.2 Application to quantum key distribution

### v3

@@ 3 Numerical comparison of the performance with other concentration inequalities
3.1 Toy model for statistical estimation
@@
$\begin{split}&{\cal A}(\gamma,\Delta)\\ &\coloneqq\bigl{\{}(p_{1},p_{2},p_{3}):p_{1},p_{2},p_{3}\geq 0,\,p_{1}+p_{2}+p_{3}=1,\\ &\hskip 75.39963ptp_{1}-\gamma p_{2}\geq m(\gamma)+\Delta\bigr{\}},\end{split}$ (156)
@@
$M_{c}(\gamma)$ $\leq m(\gamma)I,$ (158)
$m(\gamma)$ $\coloneqq\frac{1-\gamma+\sqrt{(1-\gamma)^{2}+4r\gamma}}{4},$ (159)
@@
$\begin{split}&\mathrm{Pr}\left[\frac{1}{N}\sum_{i}\left(\hat{c}_{i}-\mathbb{E}[\hat{c}_{i}\mid\hat{c}_{<i}]\right)\geq t\right]\\ &\leq\exp\bigl{[}-2Nt^{2}/(1+\gamma)^{2}\bigr{]},\end{split}$ (164)
@@
$\begin{split}&\mathrm{Pr}\left[\frac{\hat{A}}{N}-\gamma\frac{\hat{B}}{N}\geq m(\gamma)+\Delta\right]\\ &\leq\exp\bigl{[}-2N\Delta^{2}/(1+\gamma)^{2}\bigr{]}.\end{split}$ (167)
@@
$\epsilon_{\rm azuma}(\gamma,N,\Delta)=\exp\bigl{[}-2N\Delta^{2}/(1+\gamma)^{2}\bigr{]}$ (169)
@@
$\begin{split}&\mathrm{Pr}\biggl{[}\sum_{i=1}^{N}\hat{a}_{i}-\mathbb{E}[\hat{a}_{i}\mid\hat{a}_{<i},\hat{b}_{<i}]-\gamma\left(\hat{b}_{i}-\mathbb{E}[\hat{b}_{i}\mid\hat{a}_{<i},\hat{b}_{<i}]\right)\\ &\quad\geq\gamma\sqrt{N}\biggl{(}\beta^{*}-\alpha^{*}+\frac{2\alpha^{*}}{N}\sum_{i=1}^{N}\hat{b}_{i}\biggr{)}+\sqrt{\frac{N\ln(2/\epsilon)}{2}}\biggr{]}\\ &\hskip 199.16928pt\leq\epsilon,\end{split}$ (173)
@@
$\begin{split}&\Delta_{\rm kato}(\gamma,N,\epsilon)\\ &=\frac{\gamma}{\sqrt{N}}\biggl{[}\beta^{*}-\alpha^{*}\biggl{(}1-\frac{2\hat{B}}{N}\biggr{)}\biggr{]}+\sqrt{\frac{\ln(2/\epsilon)}{2N}}\end{split}$ (176)
achieves Eq. (161) in this case.
Due to the use of the union bound, the derivation of a concentration inequality of the type Eq. (162) is not carried out in the same way as the case of Azuma’s inequality. With Corollary 1 in Ref. [14] applied to $\{\hat{b}\}_{i=1}^{N}$ and $\delta$ and $\epsilon$ in Corollary 1 set to $1-2B^{*}/N$ and $\Delta^{\prime}\sqrt{N}$, respectively, we have
$\begin{split}&\mathrm{Pr}\biggl{[}\frac{1}{N}\sum_{i=1}^{N}(\mathbb{E}[\hat{b}_{i}\mid\hat{a}_{<i},\hat{b}_{<i}]-\hat{b}_{i})\geq\xi(\hat{B},B^{*},\Delta^{\prime})\Delta^{\prime}\biggr{]}\\ &\qquad\leq\exp\left[-\frac{2N\Delta^{\prime 2}}{1-(1-2\frac{B^{*}}{N}-\frac{4}{3}\Delta^{\prime})^{2}}\right],\end{split}$ (177)
where $\xi(\hat{B},B^{*},\Delta^{\prime})$ is defined as
$\xi(\hat{B},B^{*},\Delta^{\prime})\coloneqq\frac{1-(1-2\frac{\hat{B}}{N})(1-\frac{B^{*}}{N}-\frac{4}{3}\Delta^{\prime})}{1-(1-2\frac{B^{*}}{N})(1-\frac{B^{*}}{N}-\frac{4}{3}\Delta^{\prime})}.$ (178)
When $\hat{B}=B^{*}$, it is unity, i.e., $\xi(B^{*},B^{*},\Delta^{\prime})=1$. Now, given $\Delta$, choose $\Delta^{\prime}$ so that $\gamma\xi(\hat{B},B^{*},\Delta^{\prime})\Delta^{\prime}\leq\Delta$. Then, with Azuma’s inequality applied to $\{\hat{a}_{i}\}_{i=1}^{N}$, we have
$\begin{split}&\mathrm{Pr}\biggl{[}\frac{1}{N}\sum_{i=1}^{N}(\hat{a}_{i}-\mathbb{E}[\hat{a}_{i}\mid\hat{a}_{<i},\hat{b}_{<i}])\geq\Delta-\gamma\xi(\hat{B},B^{*},\Delta^{\prime})\Delta^{\prime}\biggr{]}\\ &\qquad\leq\exp\left[-2N(\Delta-\gamma\xi(\hat{B},B^{*},\Delta^{\prime})\Delta^{\prime})^{2}\right].\end{split}$ (179)
Combining Eqs. (160), (177), and (179), we have
$\mathrm{Pr}\biggl{[}\frac{\hat{A}}{N}-\gamma\frac{\hat{B}}{N}\geq m(\gamma)+\Delta\biggr{]}\leq\epsilon_{\rm kato}(\gamma,N,\Delta),$ (180)
where $\epsilon_{\rm kato}(\gamma,N,\Delta)$ is given by
$\begin{split}&\epsilon_{\rm kato}(\gamma,N,\Delta)\\ &=\min_{0\leq\gamma\xi(\hat{B},B^{*},\Delta^{\prime})\Delta^{\prime}\leq\Delta}\exp\left[-\frac{2N\Delta^{\prime 2}}{1-(1-2\frac{B^{*}}{N}-\frac{4}{3}\Delta^{\prime})^{2}}\right]\\ &\hskip 62.59596pt+\exp\left[-2N(\Delta-\gamma\xi(\hat{B},B^{*},\Delta^{\prime})\Delta^{\prime})^{2}\right].\end{split}$ (181)
@@
Now, assume that the i.i.d. state $\rho(\bm{q})^{\otimes n}$ is prepared with $\rho(\bm{q})$ giving the probability mass function $\bm{q}\in{\cal Q}$. Then, the probability that the sums of the outcomes of these measurements $\hat{A}$ and $\hat{B}$ lies in the region ${\cal A}(\gamma,\Delta_{\rm iid})$ is given from the refined version of the Sanov’s theorem [2, 3] as
$\begin{split}&\mathrm{Pr}\left[\left(\frac{\hat{A}}{N},\frac{\hat{B}}{N},1-\frac{\hat{A}}{N}-\frac{\hat{B}}{N}\right)\in{\cal A}(\gamma,\Delta_{\rm iid})\middle|\rho(\bm{q})^{\otimes n}\right]\\ &\leq\max_{\bm{p}\in{\cal A}(\gamma,\Delta_{\rm iid})}\exp[-ND(\bm{p}\|\bm{q})].\end{split}$ (183)
@@
$\begin{split}&\mathrm{Pr}\left[\left(\frac{\hat{A}}{N},\frac{\hat{B}}{N},1-\frac{\hat{A}}{N}-\frac{\hat{B}}{N}\right)\in{\cal A}(\gamma,\Delta_{\rm iid})\middle|\rho^{\otimes N}\right]\\ &\leq\max_{\bm{p}\in{\cal A}(\gamma,\Delta_{\rm iid})}\max_{\bm{q}\in{\cal Q}}\exp[-ND(\bm{p}\|\bm{q})].\end{split}$ (184)
Now, we consider a permutation-invariant state $\rho_{\rm sym}$ and apply Theorem 1 to have
@@
then we obtain a concentration inequality of the type Eq. (162). For later use, we define $\Delta_{\rm iid}(\gamma,N,\epsilon)$ and $\epsilon_{\rm iid}(\gamma,N,\Delta)$ in the same way using Eq. (184), i.e.,
$\epsilon_{\rm iid}(\gamma,N,\Delta)=\max_{\bm{p}\in{\cal A}(\gamma,\Delta)}\max_{\bm{q}\in{\cal Q}}\exp[-ND(\bm{p}\|\bm{q})].$ (187)
We finally consider applying Theorem 3 to this problem. As previously mentioned, an independent and identical measurement is performed on an unknown quantum state and the set of measurement outcomes is given by $\{(\text{I},0),(\text{I},1),(\text{I\!I},0),(\text{I\!I},1)\}$. Since we do not distinguish $(\text{I\!I},0)$ and $(\text{I\!I},1)$ as mentioned previously, we regard this measurement as three-outcome measurement and set $k=3$. For the set ${\cal Q}$ of probability mass functions that can be realized by this measurement given in Eq. (182), we consider the following set ${\cal R}(\gamma^{\prime})\supseteq{\cal Q}$:
@@
The fact that this includes ${\cal Q}$ as a subset follows from Eqs. (157) and (159). Due to the requirement of Theorem 3 that the complement ${\cal P}\setminus{\cal R}(\gamma^{\prime})$ of ${\cal R}(\gamma^{\prime})$ contains only one probability mass function that corresponds to having a deterministic outcome, the parameter $\gamma^{\prime}$ is chosen to satisfy $10\leq m(\gamma^{\prime})<$ 1. Likewise, the region ${\cal R}^{\prime}(\gamma^{\prime})$ can be defined through Eq. (58) by
@@
Figure 2: Comparison of the deviation $\Delta(\gamma,N,\epsilon)$ from (an upper bound on) the expectation value in Eqs. (156) and (161) when allowing the failure probability $\epsilon$. The parameters are chosen to be $\epsilon=10^{-30}$, $r=0.01$, and $B^{*}=0.01N$, and $\gamma$ is optimized to be $1.662$ so that an unconfirmed upper bound $\gamma B^{*}/N+m(\gamma)$ on $\hat{A}/N$ is minimized to be $0.02873$. We further assume that the random variable $\hat{B}$ is equal to the expected value $B^{*}$. The functions $\Delta_{\rm azuma}$ and $\Delta_{\rm kato}$ are defined respectively in Eqs. (168) and (176), and the functions $\Delta_{q}$, and $\Delta_{c}$ are defined implicitly through Eqs. (185) and (190). For comparison, we also plot the deviation $\Delta_{\rm iid}$ when the i.i.d. quantum state is assumed to be given, where $\Delta_{\rm iid}$ is defined implicitly through Eq. (184).
@@
We numerically demonstrate the tightness of these bounds in some cases. The first example is the case that is relevant to the finite-size analysis of the QKD. In QKD, the parameter $\epsilon$ is fixed and its value is typically chosen to be $\sim 10^{-30}$. In the QKD application, we have the expected value $B^{*}$ of $\hat{B}$ from the channel that we use for the transmission, which is the reason why Kato’s inequality [14] typically gives a tighter bound in the finite-size regime. For a model, we set $\epsilon=10^{-30}$, $r=0.01$, and $B^{*}=0.01N$. The parameter $\gamma$ in ${\cal A}(\gamma,\Delta)$ is chosen so that the expected (unconfirmed) upper bound $\gamma B^{*}/N+m(\gamma)$ on $\hat{A}/N$ in the limit $N\rightarrow\infty$ is minimized. In the choice of parameters above, we have $\gamma=1.662$ to give the expected asymptotic upper bound $0.02873$ on $\hat{A}/N$. We further assume the case in which $\hat{B}$ is exactly equal to $B^{*}$. Under these assumptions, we derive the deviation from the above asymptotic upper bound for finite $N$ allowing the failure probability $\epsilon$ with the four methods described above. Figure 2 shows the deviation $\Delta(\gamma,N,\epsilon)$ when varying the number $N$ of quantum systems to be measured. As the figure suggests, $\Delta_{c}$ defined through Eq. 190 is the smallest among four bounds in an adversarial setup (i.e., excluding $\Delta_{\rm iid}$) and with the above particular choices of parameters, except for some points that may be caused by the insufficient optimization. We can also see that $\Delta_{q}$ defined through Eq. (185) gives the comparable bound, and these two ($\Delta_{c}$ and $\Delta_{q}$) are as small as that of the i.i.d. case given by $\Delta_{\rm iid}$. They are smaller than $\Delta_{\rm kato}$ defined in Eq. (176), while $\Delta_{\rm kato}$ is still smaller than the $\Delta_{\rm azuma}$ defined in Eq. (168).
@@
Figure 3: Comparison of the failure probability $\epsilon(\gamma,N,\Delta)$ in Eq. (162) against the number $N$ of measured quantum systems for a fixed value of the deviation $\Delta$. The parameters are chosen to be $\Delta=0.01$, $r=0.01$, and $B^{*}=0.01N$, and $\gamma$ is optimized to be $1.662$ so that an unconfirmed upper bound $\gamma B^{*}/N+m(\gamma)$ on $\hat{A}/N$ is minimized to be $0.02873$. We further assume that the random variable $\hat{B}$ is equal to the expected value $B^{*}$. The functions $\epsilon_{\rm azuma}$, $\epsilon_{q}$, $\epsilon_{c}$, and $\epsilon_{\rm iid}$ are defined in Eqs. (169), (186), (191), and (187), respectively.
@@
The next example is the case when we fix the deviation $\Delta$ from the expectation value (that is, an upper bound that holds in the asymptotic limit as shown in Eq. (165)) to a small value, and see how quickly the failure rate $\epsilon(\gamma,N,\Delta)$ in Eq. (162) goes to zero. We set the deviation $\Delta=0.01$, $r=0.01$ and $B^{*}=0.01N$, and choose $\gamma=1.662$ to give the asymptotically optimal upper bound $0.02873$ on $\hat{A}/N$ as was done in the previous figure. We also consider the case in which the observed value $\hat{B}$ is exactly equal to $B^{*}$. With these conditions, we plot $\epsilon_{\rm azuma}$ in Eq. (169), $\epsilon_{\rm kato}$ in Eq. (181), $\epsilon_{q}$ in Eq. (186), $\epsilon_{c}$ in Eq. (191), and $\epsilon_{\rm iid}$ in Eq. (187) in Fig. 3. As the figure suggests, the failure probability exponentially decays, but the exponents are very different between Azuma’s inequality, Kato’s inequality, and ours. In fact, the convergence speed of our bounds is much quicker than that of Azuma’s or Kato’s inequality and as quick as that of i.i.d.. Furthermore, $\epsilon_{q}$, $\epsilon_{c}$, and $\epsilon_{\rm iid}$ appear to have almost constant gaps in the logarithmic plot. This gap exactly comes from the difference of the factors $f_{q}(N,d=2)\sim N^{\frac{3}{2}}$, $f_{c}(N,k=3)\sim N$, and $1$. We should note that there are regions in which Azuma’s and Kato’s inequality gives a better bound than ours when $N$ is small and the failure probability is large. When we alter the parameters $r$, $M$, and $\Delta$ in our model, then this region varies. However, our bound overall gives a faster exponential decay.
3.2 Application to quantum key distribution